# Attacking at non-harmonic frequencies in screaming-channel attacks

### v1

@@ Abstract
Screaming-channel attacks enable Electromagnetic (EM) Side-Channel Attacks at larger distances due to higher EM leakage energies than traditional SCAs, relaxing the requirement of close access to the victim. This attack can be mounted on devices integrating Radio Frequency (RF) modules on the same die as digital circuits, where the RF can unintentionally capture, modulate, amplify, and transmit the leakage along with legitimate signals. Leakage results from digital switching activity, so the hypothesis of previous works was that this leakage would appear at multiples of the digital clock frequency, i.e., harmonics.
This work demonstrates that compromising signals appear not only at the harmonics and that leakage at non-harmonics can be exploited for successful attacks. Indeed, the transformations undergone by the leaked signal are complex due to propagation effects through the substrate and power and ground planes, so the leakage also appears at other frequencies. We first propose two methodologies to locate frequencies that contain leakage and demonstrate that it appears at non-harmonic frequencies. Then, our experimental results show that screaming-channel attacks at non-harmonic frequencies can be as successful as at harmonics when retrieving a 16-byte AES key. As the RF spectrum is polluted by interfering signals, we run experiments and show successful attacks in a more realistic, noisy environment where harmonic frequencies are contaminated by multi-path fading and interference. These attacks at non-harmonic frequencies increase the attack surface by providing attackers with an increased number of potential frequencies where attacks can succeed.
@@ 1 Introduction
To overcome this limitation and further study the risk posed by screaming channels, this paper studies the attack’s feasibility when capturing signals at frequencies other than the harmonics of the digital processing clock. Specifically, we seek to answer the following question: is exploitable leakage also present at frequencies other than the harmonics? If this question is answered positively, attackers can have an extensive choice of potential frequencies to select from and find one that is not polluted by environmental noise during the attack. Such a property can also be an enabler to effectively extend the framework of multi-channel attacks [4], which attack by combining different side-channel sources, and combine different frequencies in the context of modulated leakage signals.
@@
  • We demonstrate successful attacks in more realistic scenarios. We apply the proposed methodologies and the insights learned when attacking at non-harmonics and build attacks in a context where most harmonics are polluted by other standard signals typically found in the spectrum.
@@ 2.1 Side-channel attacks
By capturing the leakage generated by computing devices, attackers can mount SCAs to jeopardize confidentiality and recover internal secret data. The most common methods used to build SCAs are Differential Power Analysis (DPA)[16], Correlation Power Analysis (CPA)[5], Mutual Information Analysis (MIA)[13], Template Attacks (TA)[8] and more recently Deep Learning (DL[19]. The steep peaks of the signals resulting from the switching activity of transistors in digital devices produce leakage signals whose variations correlate with that switching activity and, therefore, with the data over which the processor computes. That leakage signal is then most often found in the EM emanations [12] or the power consumption variations [18] of the victim device.
The leakage signal of interest is generated by the digital part of the system from the switching activity of the transistors occurring when data is being computed or moving through the chip. In a synchronous device, transistors switch at a pace much influenced by a unique digital clock signal. The leakage resulting from successions of transitions and non-transitions is originally modulated by this clock, which acts as an oscillator [3]. Since the clock is a square wave signal, its Fourier transform corresponds to peaks located at the harmonics $H(f)$ of its main oscillating frequency $F_{clk}$, as formalized in Equation (1), with $A_{n}$ being the respective amplitudes of each harmonic.
$\begin{split}&H(f)=\Sigma_{n=-\infty}^{\infty}A_{n}\delta(f-nF_{clk})\end{split}$ (1)
Theoretically, if the clock signal was a perfect square with a duty cycle of $50\%$, its frequency spectrum would have peaks only at the odd harmonics. However, in practice, the duty cycle is never perfect, and peaks also appear at even harmonics [1] as illustrated in Fig. 1. This is why the frequency spectrum of conventional side-channels $SC$, being the digital noise $N(f)$ modulated by the clock signal, has a peak of energy at each harmonic of the clock frequency as formalized in Equation (2).
Figure 1: Digital clock harmonics: Theoretically, since the clock signal is a square wave with a duty cycle of $50\%$, its Fourier transform should contain only the odd harmonics (green peaks). But a real clock signal does not have a duty cycle perfectly equal to $50\%$. Hence the even harmonics (blue peaks) also appear.
$SC(f)=N(f)\ast H(f)=\Sigma_{n=-\infty}^{\infty}N(f)\times A_{n}\delta(f-nF_{clk})$ (2)
A limitation of traditional SCAs to capture clean leakage signals is that attackers must be in very close proximity to the victim device for a successful attack, usually only a few millimeters away. There are, however, some specific scenarios that allow an attacker to take distance from the device [24, 10]. One scenario for gaining distance from the device, which is the focus of this paper, is the so-called screaming-channel attack [7]. In this attack, the leakage is transmitted by an RF module that sits beside the digital part of a mixed-signal chip on the same die, allowing the attacker to capture it at a distance of several meters.
@@ 2.2 Screaming-channel attacks
$ScreamC(f)=SC(f)\ast F_{RF}=\Sigma_{n=-\infty}^{\infty}N(f)\times A_{n}\delta(f-nF_{clk}-F_{RF})$ (3)
@@ 3 The attack scenario and setup
Fig. 3 shows the attack setup used for all experiments. The victim device is an nRF52832 from Nordic instrument [^1] https://www.nordicsemi.com/Software-and-tools/Development-Kits/nRF52-DK.. It contains an Arm Cortex M4 processor and an RF module. The attacker device is a USRP N210222https://www.ettus.com/all-products/un210-kit/ using an SBX daughter board that can measure a signal between $400$ MHz and $4.4$ GHz and has a bandwidth of $40$ MHz. To collect the leakage at distance, a parabolic grid antenna with a gain of $26$ dBi is used. The computer used for all our experiments has a 4-core Intel Xeon(R) CPU E3-1226 V3 @ 3.30 GHz, and $8$ GB RAM.
@@
Figure 3: Experimental setup: (2(a)) shows our wired setup for noiseless experiments and (2(b)) an attacker at 7 meters from the victim. This last one contains: 1) (blue arrow) the victim PCA10040 device running AES and transmitting a Bluetooth signal at $2.4$ GHz.; and 2) (orange arrow) the SDR device that collects the data-correlated leakage from the victim.
@@ 3.2 Trace segmentation
To extract an initial pattern, we use VT [15] as pattern extraction technique, illustrated on Fig. 5. This study shows that by knowing the precise time duration $L_{cp}$ between 2 CPs, it is possible to segment a raw trace containing leakage produced by a series of CP executions. Averaging the obtained segments returns a representative pattern of the leakage produced by that device each time it executes a CP. The study also describes a procedure to find this precise length $L_{cp}$.
Figure 5: Virtual Trigger (VT[15]: Knowing the precise duration between 2 CP executions $L_{cp}$ enables the segmentation of a raw trace. Splitting this raw trace in segments whose starts are separated by $L_{cp}$ returns segments all containing leakage of one CP execution. Averaging these segments returns a reduced-noise segment which can be used as a pattern
@@ 3.3 Time diversity
In the remaining of this paper, this reduced-noise segment is called a trace. The number $N$ is set to $10$ in the experiments where the leakage is collected through a cable (Sections 4.1.1 and 5). It is increased to $50$ for the experiments where it is collected at a distance with an antenna (Section 6). These values are significantly different from the $500$ traces used in[6] when attacking at a distance. While this brings additional difficulty in running a successful attack, it allowed us to run the experiments in a more reasonable time of 6 days instead of 5 weeks. For each experiment, the number of traces collected will be noted as $Nb_{Traces}\times N_{Time\_Diversity}$.
@@ 4 Searching for leakage at non-harmonic frequencies
Fig. 6 shows a part of the frequency spectrum at the output of the victim board while transmitting a Bluetooth signal at $2.4$ GHz. Next to the legitimate signal peak, other peaks with lower energy appear. These correspond to the first $4$ harmonics of the leakage from the digital part. They are present at frequencies equal to $2.4$ GHz plus multiples of $64$ MHz, which is the digital clock frequency. Therefore, one would intuitively assume that the leakage from the digital part transmitted by the victim board is stronger at these harmonics and hence that it would be more challenging to perform the attack at other frequencies. For this reason, previous works on screaming-channel attacks [6, 27] use harmonic frequencies for the attack. The second harmonic is typically the one used, as it is often less polluted by interfering signals, while the first harmonic is located within a frequency band used by signals like Bluetooth or WiFi. Nevertheless, it can be seen on the spectrum that some variations in energy are also present between these harmonics, suggesting that leakage could potentially be distributed continuously along the spectrum.
@@
The first question we want to answer is: is the leakage from the digital part only present at the harmonic frequencies, or does it also appear at other frequencies? To answer this question, we propose two methodologies to investigate at which frequencies the leakage exists in the spectrum. The first consists in running a t-test at each tested frequency. This test is commonly used by the side-channel community to determine whether the internal data computed by the CPs has an impact on the leakage. Then we propose a second method that reduces the implementation complexity while giving similar results. It is an adaptation of the method used in [15] to extract the CP leakage pattern; we refer to this second method as pattern detection. It consists in analyzing whether or not a signal collected at a given frequency is good enough to extract a CP pattern; if yes, this means that leakage is present. Fig. 7 illustrates the difference between the two methods, described in the remainder of this section. Compared with the first methodology, the second method removes the most time-consuming phase, which is the collection of (500) synchronized traces needed for the t-test.
Figure 7: Comparison of the steps of the two methods. Both first segment a raw trace using VT. Then, the first builds a pattern used for the collection phase and performs a t-test with the $N$ collected traces (in the presented experiment, $N=500$). Whereas the second uses the segments returned by VT to test if the raw trace actually contains leakage.
@@ 4.1 Leakage localization using t-test method
where û$_{i}$ represents the average of the samples belonging to the $i_{th}$ set, $\sigma_{i}^{2}$ is their variance, and $N_{i}$ the number of samples in this set. Thus, for the t-test to work, it is necessary to synchronize the collection of the leakage with the execution of CPs perfectly to know which samples correspond to which time point from one leakage collection to another. The leakage collection method in section 3 is used for this. For pattern recognition, a different pattern must be extracted for each frequency, as the shape of the CP leakage differs among frequencies. The VT [15] technique is used to automate pattern extraction.
In the experiment we collect $500$ traces at each frequency using the wired setup from Fig 2(a). This methodology is used over the frequency range 1.4GHz to 3.4GHz ($2.4$ +/- $1$ GHz), with a resolution of 1 MHz. These parameters are application-dependent and can be adjusted according to each particular use case in order to test a wider frequency band or to change the resolution.
@@ 4.1.1 Experimental results
Fig. 8 shows the results of the t-test. At each frequency, the maximum absolute value of the t-test is kept as a score. Frequencies corresponding to the harmonics are highlighted in blue. If the leakage had only been present at these frequencies, the score would have been higher than this threshold only at these locations, but it can be observed that the score is also above 4.5 at other frequencies, suggesting that leakage is also present at non-harmonics. In fact, even if the highest peaks are at the first $2$ harmonics, the score is very high around the first $3$ harmonics and is above the threshold over a band of more than $500$ MHz of width (almost until the $6th$ harmonic). This is true on both the right and left sides of the spectrum.
@@ 5.1 The attack and score
In many cases, some bytes are incorrectly guessed, but it is still possible to brute-force the correct key. A brute-force approach [22] consists in testing the ranked keys from the most probable, according to the probabilities computed during the attack, until finding the right one. The number of keys tested by the brute-force algorithm before reaching the correct one is the Key Rank and is representative of the complexity to recover the key. The lower the key rank is, the better the attack performs, as the brute-force attack needs less time to reach the correct key. When the key rank is lower than $2^{32}$, it takes about $5$ minutes on the experimental computer to brute-force the key. When lower than $2^{35}$, the brute-force takes about $1$ hour. In the remainder of this paper, this key rank is kept as the criteria to evaluate the efficiency of an attack as in previous works [6].
@@ 6.3 Attacking at a further distance
As screaming-channel attacks try to enable attacks where attackers are as far as possible from the victim, we run a new round of attack experiments with the antenna put at a distance of $7$ meters. This time 5050505066650505050 is the minimal number usually considered by the side-channel community for statistically meaningful results attacks are performed under the exact same conditions at each frequency. To keep the experiments feasible in a reasonable time, we reduced the number of tested frequencies. The frequencies selected for this attack at $7$ meters are chosen among the ones where the attack performed best in the 2-meters scenario using $750$ traces. The results of the attacks at $7$ meters are shown in Fig. 11(d). In this case there is only one harmonic that still gets a key rank lower than $2^{35}$, while most of the selected non-harmonics still work. Among the $9$ which were selected, $5$ have a lower rank than $2^{32}$ and $2$ than $2^{35}$.
@@
Figure 12: Experimental results: Key Rank (the lower, the better). Underlying colors per frequency indicate key ranks of: green <=<=< = to $2^{32}$ (approx. $5$ mins brute-force), blue <=<=< = to $2^{35}$ (approx. $1$ hour brute-fore), red $>$ to $2^{35}$. For (11(c)) and (11(d)), the graph shows: (1) a central box with the distribution of scores from the first quartile to the third; (2) whiskers with the extension of the remaining scores by 1.5×1.5$\times$1.5 × the inter-quartile range (equal to the difference between the third and first quartile); and (3) a horizontal line with the median score of the attacks.

### v2

@@ Abstract
Screaming-channel attacks enable Electromagnetic (EM) Side-Channel Attacks at larger distances due to higher EM leakage energies than traditional SCAs, relaxing the requirement of close access to the victim. This attack can be mounted on devices integrating Radio Frequency (RF) modules on the same die as digital circuits, where the RF can unintentionally capture, modulate, amplify, and transmit the leakage along with legitimate signals. Leakage results from digital switching activity, so previous works hypothesized that this leakage would appear at multiples of the digital clock frequency, i.e., harmonics.
This work demonstrates that compromising signals appear not only at the harmonics and that leakage at non-harmonics can be exploited for successful attacks. Indeed, the transformations undergone by the leaked signal are complex due to propagation effects through the substrate and power and ground planes, so the leakage also appears at other frequencies. We first propose two methodologies to locate frequencies that contain leakage and demonstrate that it appears at non-harmonic frequencies. Then, our experimental results show that screaming-channel attacks at non-harmonic frequencies can be as successful as at harmonics when retrieving a 16-byte AES key. As the RF spectrum is polluted by interfering signals, we run experiments and show successful attacks in a more realistic, noisy environment where harmonic frequencies are contaminated by multi-path fading and interference. These attacks at non-harmonic frequencies increase the attack surface by providing attackers with more potential frequencies where attacks can succeed.
@@ 1 Introduction
To overcome this limitation and further study the risk posed by screaming channels, this paper studies the attack’s feasibility when capturing signals at frequencies other than the harmonics of the digital processing clock. Specifically, we seek to answer the following questions: is exploitable leakage also present at frequencies other than the harmonics? In case it is, is the difficulty for a successful attack higher at non-harmonics than at harmonics frequencies? If the first question is answered positively, attackers can have an extensive choice of potential frequencies to select from and find one not polluted by environmental noise during the attack. Such a property can also be an enabler to effectively extend the framework of multi-channel attacks [4], which attack by combining different side-channel sources and different frequencies in the context of modulated leakage signals.
@@
  • We demonstrate successful attacks in more realistic scenarios. We apply the proposed methodologies, and the insights learned when attacking at non-harmonics and build attacks in a context where most harmonics are polluted by other standard signals typically found in the spectrum.
@@ 2.1 Side-channel attacks
By capturing the leakage generated by computing devices, attackers can mount SCAs to jeopardize confidentiality and recover internal secret data. The most common methods used to build SCAs are Differential Power Analysis (DPA)[16], Correlation Power Analysis (CPA)[5], Mutual Information Analysis (MIA)[13], Template Attacks (TA)[8] and more recently Deep Learning (DL[19]. The leakage signal of interest is generated by the digital part of the system from the switching activity of the transistors occurring when data is being computed or moved through the chip. The steep peaks of the signals resulting from the switching activity of transistors in digital devices produce leakage signals whose variations correlate with that switching activity and, therefore, with the data over which the processor computes. That leakage signal is most often found in the EM emanations [12] or the power consumption variations [18] of the victim device.
In a synchronous device, transistors switch at the pace of the digital clock. As a result, the leakage resulting from successions of transitions and non-transitions has a period equal to one clock cycle. The Fourier Transform of a signal having a period $T$ corresponds to peaks at each harmonic of the frequency $F$ = $1/T$ [1] as formalized in Equation (1), with $A_{n}$ being the respective amplitudes of each harmonic. Then, theoretically, the leakage should appear at harmonics of $F_{clk}$ = $1/T_{clk}$ as formalized in Equation (2) and illustrated in Fig. 1. This corresponds to what is empirically observed in works on side-channel attacks where the leakage is found at harmonics of the clock frequency [3].
$\begin{split}&H(f)=\Sigma_{n=-\infty}^{\infty}A_{n}\delta(f-n/T)\end{split}$ (1)
Figure 1: Leakage presence over the spectrum: The Fourier transform of a signal having a period Tclk consists of peaks at both odd harmonics (green peaks) and even harmonics (blue peaks) of frequency $F_{clk}$ = $1/T_{clk}$.
$SC(f)=\Sigma_{n=-\infty}^{\infty}A_{n}\delta(f-nF_{clk})$ (2)
A limitation of traditional SCAs to capture clean leakage signals is that attackers must be in very close proximity to the victim device for a successful attack, usually only a few millimeters away. However, some specific scenarios allow attackers to take distance from the device [24, 10]. One scenario for gaining distance from the device, which is the focus of this paper, is the so-called screaming-channel attack [7]. In this attack, the leakage is transmitted by an RF module that sits beside the digital part of a mixed-signal chip on the same die, allowing the attacker to capture it at a distance of several meters.
@@ 2.2 Screaming-channel attacks
$ScreamC(f)=SC(f)\ast F_{RF}=\Sigma_{n=-\infty}^{\infty}A_{n}\delta(f-nF_{clk}-F_{RF})$ (3)
@@ 3 The attack scenario and setup
Fig. 3 shows the attack setup used for all experiments. The victim device is an nRF52832 from Nordic instrument [^1] https://www.nordicsemi.com/Software-and-tools/Development-Kits/nRF52-DK.. It contains an Arm Cortex M4 processor and an RF module. The attacker device is a USRP N210222https://www.ettus.com/all-products/un210-kit/ using an SBX daughter board that can measure a signal between $400$ MHz and $4.4$ GHz and has a bandwidth of $40$ MHz. To collect the leakage at a distance, a parabolic grid antenna with a gain of $26$ dBi is used. The computer used for all our experiments has a 4-core Intel Xeon(R) CPU E3-1226 V3 @ 3.30 GHz, and $8$ GB RAM.
@@
Figure 3: Experimental setup: (3(a)) shows our wired setup for noiseless experiments and (3(b)) an attacker at 7 meters from the victim. This last one contains: 1) (blue arrow) the victim PCA10040 device running AES and transmitting a Bluetooth signal at $2.4$ GHz.; and 2) (orange arrow) the SDR device that collects the data-correlated leakage from the victim.
@@ 3.2 Trace segmentation
To extract an initial pattern, we use VT [15] as the pattern extraction technique, illustrated in Fig. 5. This study shows that by knowing the precise time duration $L_{cp}$ between 2 CPs, it is possible to segment a raw trace containing leakage produced by a series of CP executions. Averaging the obtained segments returns a representative pattern of the leakage produced by that device each time it executes a CP. The study also describes a procedure to find this precise length $L_{cp}$.
Figure 5: Virtual Trigger (VT[15]: Knowing the precise duration between 2 CP executions $L_{cp}$ enables the segmentation of a raw trace. Splitting this raw trace in segments whose starts are separated by $L_{cp}$ returns segments all containing leakage of one CP execution. Averaging these segments returns a reduced-noise segment, which can be used as a pattern
@@ 3.3 Time diversity
In the rest of this paper, this reduced-noise segment is called a trace. The number $N$ is set to $10$ in the experiments where the leakage is collected through a cable (Sections 4.1.1 and 5). It is increased to $50$ for the experiments where it is collected at a distance with an antenna (Section 6). These values significantly differ from the $500$ traces used in[6] when attacking at a distance. While this brings additional difficulty in running a successful attack, it allowed us to run the experiments in a more reasonable time of 6 days instead of 5 weeks. For each experiment, the number of traces collected will be noted as $Nb_{Traces}\times N_{Time\_Diversity}$.
@@ 4 Searching for leakage at non-harmonic frequencies
Fig. 6 shows a part of the frequency spectrum at the output of the victim board while transmitting a Bluetooth signal at $2.4$ GHz. Next to the legitimate signal peak, other peaks with lower energy appear. These correspond to the first $4$ harmonics of the leakage from the digital part. They are present at frequencies equal to $2.4$ GHz plus multiples of $64$ MHz, which is the digital clock frequency. Therefore, one would intuitively assume that the leakage from the digital part transmitted by the victim board is stronger at these harmonics and, hence, that it would be more challenging to perform the attack at other frequencies. For this reason, previous works on screaming-channel attacks [6, 27] use harmonic frequencies for the attack. Typically, the second harmonic is used, as it is often less polluted by interfering signals. In contrast, the first harmonic is located within a frequency band typically used by signals like Bluetooth or WiFi. Nevertheless, it can be seen on the spectrum that some variations in energy are also present between these harmonics, suggesting that leakage could potentially be distributed continuously along the spectrum.
@@
The first question we want to answer is: is the leakage from the digital part only present at the harmonic frequencies, or does it also appear at other frequencies? To answer this question, we propose two methodologies to investigate at which frequencies the leakage exists in the spectrum. The first involves running a t-test at each tested frequency. The side-channel community commonly uses this test to determine whether the internal data computed by the CPs impacts the leakage. Then, we propose a second method that reduces the implementation complexity while giving similar results. It is an adaptation of the method used in [15] to extract the CP leakage pattern; we refer to this second method as pattern detection. It consists of analyzing whether a signal collected at a given frequency is good enough to extract a CP pattern; if yes, this means that leakage is present. Fig. 7 illustrates the difference between the two methods, described in the remainder of this section. Compared with the first methodology, the second method removes the most time-consuming phase: the collection of (500) synchronized traces needed for the t-test.
Figure 7: Comparison of the steps of the two methods. Both first segment a raw trace using VT. Then, the first builds a pattern used for the collection phase and performs a t-test with the $N$ collected traces (in the presented experiment, $N=500$). The second uses the segments returned by VT to test if the raw trace contains leakage.
@@ 4.1 Leakage localization using t-test method
where û$_{i}$ represents the average of the samples belonging to the $i_{th}$ set, $\sigma_{i}^{2}$ is their variance, and $N_{i}$ the number of samples in this set. Thus, for the t-test to work, it is necessary to synchronize the leakage collection with the execution of CPs perfectly to know which samples correspond to which time point from one leakage collection to another. The leakage collection method in section 3 is used for this. For pattern recognition, a different pattern must be extracted for each frequency, as the shape of the CP leakage differs among frequencies. The VT [15] technique is used to automate pattern extraction.
In the experiment we collect $500$ traces at each frequency using the wired setup from Fig 3(a). This methodology is used over the frequency range 1.4GHz to 3.4GHz ($2.4$ +/- $1$ GHz), with a resolution of 1 MHz. These parameters are application-dependent and can be adjusted according to each particular use case to test a wider frequency band or to change the resolution.
@@ 4.1.1 Experimental results
Fig. 8 shows the results of the t-test. At each frequency, the maximum absolute value of the t-test is kept as a score. Frequencies corresponding to the harmonics are highlighted in blue. If the leakage had only been present at these frequencies, the score would have been higher than this threshold only at these locations. Still, it can be observed that the score is also above 4.5 at other frequencies, suggesting that leakage is also present at non-harmonics. In fact, even if the highest peaks are at the first $2$ harmonics, the score is very high around the first $3$ harmonics and is above the threshold over a band of more than $500$ MHz of width (almost until the $6th$ harmonic). This is true on both the right and left sides of the spectrum.
@@ 5.1 The attack and score
In many cases, some bytes are incorrectly guessed, but it is still possible to brute-force the correct key. A brute-force approach [22] consists in testing the ranked keys from the most probable, according to the probabilities computed during the attack, until finding the right one. The number of keys tested by the brute-force algorithm before reaching the correct one is the Key Rank and is representative of the complexity of recovering the key. The lower the key rank is, the better the attack performs, as the brute-force attack needs less time to reach the correct key. When the key rank is lower than $2^{32}$, it takes about $5$ minutes on the experimental computer to brute-force the key. When lower than $2^{35}$, the brute-force takes about $1$ hour. In the remainder of this paper, this key rank is kept as the criteria to evaluate the efficiency of an attack as in previous works [6].
@@ 6.3 Attacking at a further distance
As screaming-channel attacks try to enable attacks where attackers are as far as possible from the victim, we run a new round of attack experiments with the antenna put at a distance of $7$ meters. This time 5050505066650505050 is the minimal number usually considered by the side-channel community for statistically meaningful results attacks are performed under the exact same conditions at each frequency. To keep the experiments feasible in a reasonable time, we reduced the number of tested frequencies. The frequencies selected for this attack at $7$ meters are chosen among the ones where the attack performed best in the 2-meters scenario using $750$ traces. The results of the attacks at $7$ meters are shown in Fig. 12(d). In this case, there is only one harmonic that still gets a key rank lower than $2^{35}$, while most of the selected non-harmonics still work. Among the $9$ which were selected, $5$ have a lower rank than $2^{32}$ and $2$ than $2^{35}$.
@@
Figure 12: Experimental results: Key Rank (the lower, the better). Underlying colors per frequency indicate key ranks of: green <=<=< = to $2^{32}$ (approx. $5$ mins brute-force), blue <=<=< = to $2^{35}$ (approx. $1$ hour brute-fore), red $>$ to $2^{35}$. For (12(c)) and (12(d)), the graph shows: (1) a central box with the distribution of scores from the first quartile to the third; (2) whiskers with the extension of the remaining scores by 1.5×1.5$\times$1.5 × the inter-quartile range (equal to the difference between the third and first quartile); and (3) a horizontal line with the median score of the attacks.